OpenStudy (anonymous):

Colorado outfielder Matt Holliday had a batting average of .340 in 2007. a) What is the probability that Holliday would go 0-for-4 (zero hits in four at-bats) in a game in 2007? b) If Holliday had 650 at bats in 2007, what would be his expected number of hits?

OpenStudy (anonymous):

well this time we're going to use the hypergeometric distribution for part a: HG(x, t)[4] with H(HollidayAtBat) = 0.340 * 2007 P(At-Bat) = 1.0 - P(AllMiss) = 0.1252 so 0.1252 is final answer Part b just multiply 650*2007*0.340 / (223*3^2) = 221

OpenStudy (anonymous):

Wait sorry I did part A wrong, I throught it said 5 at-bats

OpenStudy (anonymous):

It should be 0.19

OpenStudy (anonymous):

ok thnx